X²-y²-c²+2yc÷((x+y-c)/(x+y+c))
1
Expert's answer
2020-08-02T16:06:25-0400

"simplify:\\
x^2-y^2-c^2+\frac{2yc}{\frac{x+y-c}{x+y+c}}=\\
=x^2-y^2-c^2+2yc\cdot \frac{x+y+c}{x+y-c}=\\
=x^2-y^2-c^2+2yc\cdot \frac{x+y-c+2c}{x+y-c}=\\
=x^2-y^2-c^2+2yc+\frac{4yc^2}{x+y-c}=\\
=x^2-(y-c)^2+\frac{4yc^2}{x+y-c}=\\
=(x-y+c)(x+y-c)+\frac{4yc^2}{x+y-c};"